# On the Moduli Space of Discrete Morse Functions

### v1

@@ Abstract.
In this work, we give a combinatorial-geometric model for moduli spaces of discrete Morse functions on CW complexes. We investigate properties of moduli spaces of discrete Morse functions and their relationship to the spaces of smooth Morse functions on manifolds, as well as spaces of discrete Morse matchings on CW complexes. Moreover, we provide a similar model for the moduli space of merge trees and realise the induced merge tree and the induced barcode as continuous maps between moduli spaces.
@@ 1. Introduction
Discrete Morse theory, originally introduced by Forman [For98], is a powerful framework for investigating the simple homotopy type and other properties of CW complexes. It has shown potential to be useful for the study of persistent invariants of filtered CW complexes as they appear in topological data analysis. Other than that, discrete Morse theory has numerous applications in both, pure and applied mathematics.
Moduli spaces are most prevalent in algebraic geometry but have also been studied in other fields of mathematics, mostly other subfields of geometry and topology. The general idea of moduli spaces is to provide a topological space $M$, preferably together with a convenient way to parametrize $M$, such that points of $M$ correspond to certain mathematical objects one wants to investigate and geometric features of $M$ reflect relevant properties of said objects under investigations.
The framework we choose for the moduli-space-approach to discrete Morse functions is hyperplane arrangements in vector spaces of discrete functions. We observe that Formans definition of discrete Morse functions Definition 3.2 gives rise to the Definition 3.9 of the Morse arrangement $\mathcal{A}(X)$, a hyperplane arrangement that subdivides the vector space of discrete functions $\mathbb{R}^{X}$ on a CW complex $X$ into regions, some of which belong to Forman-equivalence classes of discrete Morse functions, and others do not contain discrete Morse functions at all. This way, the introduced moduli space of discrete Morse functions allows a parametrization indexed over the face poset $D(X)$, which helps us to investigate properties of discrete Morse functions using the combinatorial properties of the Morse arrangement $\mathcal{A}$.
In this work, we want to investigate discrete Morse theory and its connection to certain other concepts in mathematics from the viewpoint of moduli spaces.
The concepts we want to relate the moduli spaces of discrete Morse funcitons to are
@@
    The moduli spaces of smooth Morse functions on manifolds.
@@
    The moduli spaces of discrete Morse matchings on regular CW complexes.
@@
    The moduli space of merge trees.
  4. (4)
    The moduli space of barcodes.
The moduli spaces of Morse functions on manifolds have been introduced in [Cer70] and further investigated in [HW73] as subspaces of the spaces of smooth functions $C^{\infty}(M)$ on compact smooth manifolds $M$, governed by a stratification given by a certain notion of regularity for critical points and critical values. The motivation to investigate Morse functions in this way was to give a solid framework for the investigation of the pseudo-isotopoy versus isotopy question and its implications for the h-cobordism theorem. Cerf already found a map that relates certain subspaces of the space of Morse functions to spaces of discrete Morse functions Proposition 4.2 decades before discrete Morse functions had been invented by Forman. We extend this map to connected components:
@@
It follows from Theorem 4.8 that Cerf’s map from Proposition 4.2 extends to a map $\eta\colon\mathcal{M}(\mathcal{N})\rightarrow\mathcal{M}_{\mathcal{N}}$, where $\mathcal{M}_{\mathcal{N}}$ denotes the CW decomposition of $M$ induced by any Morse function in the connected component $\mathcal{N}$ of $M$. Moreover, Cerf’s proof of Proposition 4.2 shows that this instance of $\eta$ is a topological submersion and compatible with the respective stratifications, too.
The moduli space of discrete Morse matchings has been introduced in [CJ05] and further investigated in [CM17] and [LS21]. We show that the moduli space of discrete Morse matchings is also canonically associated to the Morse arrangement:
@@
For the moduli space of merge trees, we follow a similar approach as for the moduli space of discrete Morse functions. We model merge trees as maps from combinatorial merge trees to the real numbers and topologize using a mixture of the euclidean distance and a kind of edit distance. Definition 5.15.
@@
The euclidean edit distance $\operatorname{d}$ from Definition 5.15 is a pseudo metric on the moduli space of merge trees $Mer$ and a metric on the moduli space of strict merge trees $Mer_{<}$ and the moduli space of well-branched merge trees $Mer_{wb}$.
@@
Let $X$ be a regular CW complex. Then the map $M\colon\mathcal{M}(X)\rightarrow Mer$ from Definition 5.8 that maps a discrete Morse function to its induced merge tree is continuous.
For the relationship to the moduli space of barcodes, we use a similar approach [^1] One can identify the space of barcodes over a given combinatorial barcode in this work with the space $\mathbb{R}^{2n}$ in [BG23, Section 4.1]. as in [BG23] for the space of barcodes and a construction for the barcode induced by a merge tree similar to the one in [${}^{+}$24]. We obtain the following result:
@@
The map $B\colon Mer_{wb}\rightarrow Bar$ from Definition 5.28 is continuous.
@@ Acknowledgements
The author would like to thank Max Planck Institute for Mathematics for the great scientific environment in which this project was conducted. Moreover, the author would like to thank Florian Kranhold and Paul Mücksch for helpful discussions about the project. Most notably the author thanks his advisor, Viktoriya Ozornova, for her advice, many helpful discussions, and detailed feedback at multiple occasions.
@@ 2. Hyperplane arrangements and their associated spaces
We review some well established notions from combinatorial geometry and topology, which will prove to be useful for our endeavor. The standard references we refer to are [${}^{+}$] and [AM17]. While most notions are standard in the literature, we adapt and extend them to the setting of this article.
Definition 2.1 (compare [${}^{+}$] and [AM17]).
@@
We call the connected components of the complement $V\setminus\mathcal{A}$ the (open) regions of the arrangement $\mathcal{A}$. Moreover, we call intersections of half-spaces associated to $\mathcal{A}$, with at least one half-space per hyperplane chosen, faces of the hyperplane arrangement. For oriented hyperplane arrangements, each of these faces $\sigma$ is associated with a sign vector $w\in\{+,-,0\}^{\mathcal{A}}$, where $+$ means that $\sigma$ lies on the positive side of the corresponding hyperplane, -- means that $\sigma$ lies on the negative side of the corresponding hyperplane, and 00 means that $\sigma$ lies inside corresponding hyperplane. The regions correspond to sign vectors with no entry 0. The face poset $\Sigma(\mathcal{A})$ is the set of faces of $\mathcal{A}$ ordered by inclusion.
While hyperplane arrangements can be defined in arbitrary vector spaces, it is often useful to use coordinate spaces, in particular partially ordered coordinate spaces. This allows us to use equations in the ordered coordinates in order to specify the arrangements we consider.
@@
A partially ordered coordinate space is a real vector space $V$ together with a choice of a basis which is indexed over some partially ordered set $P$.
For convenience, we sometimes choose a specific model for partially ordered coordinate spaces:
@@
The category of partially ordered coordinate spaces $poVect$ consists of
@@ 1. Objects:
    vector spaces $\operatorname{Map}(P,\mathbb{R})$ for posets $P$ together with the standard basis induced by $P$, and
@@ 2. Morphisms:
In this work, we mostly consider partially ordered coordinated spaces where the bases are indexed over either face posets of CW complexes or posets associated to merge trees or barcodes. We can characterize these three frameworks as certain types of posets, which makes partially ordered coordinate vector spaces a unifying framework for them:
@@
A poset $P$ is of regular CW type if it is isomorphic to the face poset $D(X)$ of a regular CW complex $X$.
@@
    each non-minimal element $a\in P$ has at least two elements $b\neq b^{\prime}$ such that $b<a$ and $b^{\prime}<a$.
@@
In order to see that Definition 2.4 models merge trees, we remark (2) and (3) model arbitrary rooted trees: the chains in (3) correspond to shortest paths between the root and other nodes of the tree. The feature that inner nodes of merge trees correspond to mergers of connected components, i.e. each inner node has at least two children, is reflected in (4).
@@
We denote by $\mathcal{H}_{n}$ the hyperplane arrangement in $\mathbb{R}^{n}$ given by the equations $x_{i}=x_{j}$ for all $i\neq j$. We call $\mathcal{H}_{n}$ the braid arrangement $\mathcal{H}_{n}$ (see [${}^{+}$, Example 2.3.3]).
@@
If $\mathcal{A}$ is central, then $\mathcal{L}(\mathcal{A})$ is a geometric lattice of rank $r(\mathcal{A})$. Since all arrangements in this work are central, we also refer to $\mathcal{L}(\mathcal{A})$ as the intersection lattice of $\mathcal{A}$. It is straightforward to see that $\mathcal{L}(\mathcal{A})$ always has the structure of a face poset of a simplicial complex, whereas $\Pi(\mathcal{A})$ does not necessarily satisfy this property. In particular, $\Pi(\mathcal{A})$ is never the face poset of a simplicial complex if $\mathcal{A}$ is central and $\lvert\mathcal{A}\rvert>1$.
On the other hand, $\Sigma(X)$ always has the structure of the face poset of a regular CW complex (see [AM17, 1.1.8]).
@@
We call a hyperplane arrangement simplicial if $\Sigma(\mathcal{A})\setminus\{\mathcal{Z}(\mathcal{A})\}$ has the structure of a face poset of a pure [^2] A simplicial complex is called pure if all its facets have the same dimension $d$. simplicial complex.
In fact, $\Sigma(\mathcal{A})\setminus\{\mathcal{Z}(\mathcal{A})\}$ always gives a regular CW decomposition of the sphere in $\mathcal{Z}^{\perp}$.
3. The moduli space of discrete Morse functions
Our goal is to present a framework which allows us to analyze geometric properties of and homotopies between discrete Morse functions on (regular) CW complexes. It turns out that manipulations of discrete Morse functions, like cancellations of critical cells and reordering critical cells by swapping their critical values can be geometrically realized by homotopies between discrete Morse functions, i.e. paths in the space of discrete Morse functions.
In order to define the space of discrete Morse functions on a (regular) CW complex, we consider the space of all discrete functions, which the space of discrete Morse functions will turn out to be a subspace of. As common when working with discrete Morse theory, we adopt a slight abuse of notation by not distinguishing between a CW complex and its set of cells.
@@ 3. The moduli space of discrete Morse functions
Let $X$ be a CW complex. We call the space $\mathbb{R}^{X}\coloneqq\prod\limits_{\sigma\in X}\mathbb{R_{\sigma}}$ the space of discrete functions on $X$.
Here, we notice that the coordinates of the space of discrete functions are partially ordered rather than linearly. This partial order will help us to investigate which of the discrete functions are discrete Morse functions.
@@
Let $X$ be a CW complex. A function $f\colon X\rightarrow\mathbb{R}$ is a discrete Morse function if for every $p$-dimensional cell $\alpha^{(p)}\in X$ the function $f$ is constant on the interior $\operatorname{int}(\alpha^{(p)})$ and
@@
    #$\#\{\beta^{(p+1)}\supset\alpha|f(\beta)\leq f(\alpha)\}\leq 1$, and
@@
    #$\#\{\gamma^{(p-1)}\subset\alpha|f(\gamma)\geq f(\alpha)\}\leq 1$.
@@
We include the general version for non-regular CW complexes mostly for the comparison to the smooth case in Section 4. For obvious reasons, we can drop condition (3) if $X$ is regular. It is highly inconvenient to work with discrete Morse theory on non-regular CW complexes because in that case the face poset does not contain all information of the homotopy type at hand. Moreover, we make the definition for the non-regular case a bit stricter than Forman in order to simplify the definition of the induced gradient field, i.e. Morse functions are automatically strictly monotone along face inclusions of codimension higher than one.
We recall that by [For98, Lemma 2.5] for each cell $\alpha\in X$ at most one of the inequalities (1) and (2) from Definition 3.2 may actually be an equality, whereas the other one has to be a strict inequality. Therefore, for each cell $\alpha\in X$ such that (1)/(2) is an equality there is exactly one cell $\beta\supset\alpha$/$\gamma\subset\alpha$ such that $f(\beta)\leq f(\alpha)$/$f(\alpha)\leq f(\gamma)$. Hence, any discrete Morse function $f\colon X\rightarrow\mathbb{R}$ induces a partial matching $\nabla f$ on $X$ by matching $\alpha$ with $u(\alpha)\coloneqq\beta$/$d(\alpha)\coloneqq\gamma$. The unmatched cells are called critical. To be precise, $\nabla f$ is a partial matching on the Hasse diagram $D(X)$ of the face poset of $X$:
@@
Let $X$ be a CW complex. The Hasse diagram of the face poset of $X$ is the directed graph $D(X)$ that has vertices $V(D(X))\coloneqq X$, that is the cells of $X$, and an oriented edge $(\tau,\sigma)$ whenever $\sigma\subset\tau$ is a face relation such that there exists no $\theta\in X$ with $\sigma\subset\theta\subset\tau$ ($\sigma$ is covered by $\tau$).
@@
In case of a regular CW complex $X$, the directed edges correspond to face relations of codimension one. For non-regular CW complexes, edges might correspond to face relations of higher codimension, which makes finding elementary collapses via matchings harder. Furthermore, for non-regular CW complexes, the gradient paths needed for the boundary map of the Morse–Smale complex are harder to spot inside the modified Hasse diagram, since they are then not just zig-zags between any matched edges and unmatched egdges.
@@
Let $X$ be an abstract regular CW complex. A function $f\colon X\rightarrow\mathbb{R}$ is a Morse–Benedetti function if it satisfies:
  • Monotonicity
    For cells $\sigma\subset\tau$ we have $f(\sigma)\leq f(\tau)$.
  • Semi-injectivity
    $\lvert f^{-1}(\{z\})\rvert\leq 2$ for all $z\in\mathbb{R}$.
  • Generacity
For some purposes, it is useful to weaken Benedettis definition:
@@
Let $X$ be an abstract regular CW complex. We call a function $f\colon X\rightarrow\mathbb{R}$ a weak Morse–Benedetti function if it is a discrete Morse function and satisfies semi-injectivity and generacy in the sense of Definition 3.6.

### v2

# The space of Discrete Morse Functions

@@ Abstract.
In this work, we introduce a combinatorial-geometric model for the space of discrete Morse functions on any CW complex $X$. We relate this version of a space of discrete Morse functions to the space of cellular filtrations of $X$ and discuss its relationship to various concepts such as smooth Morse theory, Cerf theory, complexes of discrete Morse matchings, and induced merge trees and barcodes.
@@ 1. Introduction
Discrete Morse theory, originally introduced by Forman [FOR98], is a powerful framework for investigating the simple homotopy type and other properties of CW complexes. It has proved to be useful for the study of persistent invariants of filtered CW complexes as they appear in topological data analysis. Other than that, discrete Morse theory has numerous applications in both, pure and applied mathematics.
The idea of studying combinatorial and geometric properties of the collection of mathematical objects of a certain kind originates in the concept of moduli spaces as they are used in topology and algebraic geometry. The general idea of moduli spaces is to provide a topological space $\mathcal{M}$, together with a convenient way to parametrize $\mathcal{M}$, such that points of $\mathcal{M}$ correspond to equivalence classes of certain mathematical objects one wants to investigate, and geometric features of $\mathcal{M}$ reflect relevant properties of said objects under investigations.
We choose the framework of hyperplane arrangements in vector spaces of discrete functions to study the space of discrete Morse functions in a way that allows the study of various properties and equivalence classes of filtrations by combinatorial-geometric methods.
We describe how Formans definition of discrete Morse functions Definition 3.2 gives rise to the Definition 3.9 of the Morse arrangement $\mathcal{A}(X)$, a hyperplane arrangement that subdivides the vector space of discrete functions $\mathbb{R}^{X}$ on a CW complex $X$ into regions, some of which represent Forman-equivalence classes of discrete Morse functions, and others do not contain discrete Morse functions at all. This way, the introduced space of discrete Morse functions allows a parametrization indexed over the face poset $F(X)$, which helps us to investigate the structure of discrete Morse functions using the combinatorial properties of the Morse arrangement $\mathcal{A}(X)$.
In this work, we investigate the space of discrete Morse functions on a given CW complex $X$ and its connection to certain other concepts in mathematics from a combinatorial-geometric viewpoint.
We provide constructions of the following spaces out of spaces of discrete Morse functions:
@@
    The spaces of discrete Morse matchings on regular CW complexes.
@@
    The space of merge trees.
@@
    The space of barcodes.
Similar work has been done in [CCF+20], [LT22], and [CMW20]. Whereas [CCF+20] explores different notions of equivalences of smooth Morse functions on the sphere and the corresponding moduli spaces, both [LT22] and [CMW20] investigate fibers of two different instances of persistence maps.
The spaces of Morse functions on manifolds have been introduced in [CER70] and further investigated in [HW73] as subspaces of the spaces of smooth functions $C^{\infty}(M)$ on compact smooth manifolds $M$, governed by a stratification given by a certain notion of regularity for critical points and critical values. The motivation to investigate Morse functions in that way was to give a solid framework for the investigation of the pseudo-isotopy versus isotopy question and its implications for the h-cobordism theorem. Cerf already found a map that relates certain subspaces of the space of Morse functions to spaces of discrete Morse functions Proposition 4.2 decades before discrete Morse functions had been invented by Forman. We extend this map to path components:
@@
It follows from Theorem 4.7 that Cerf’s map from Proposition 4.2 extends to a map $\eta\colon\mathcal{N}\rightarrow\mathcal{M}(M_{\mathcal{N}})$, where $M_{\mathcal{N}}$ denotes the CW decomposition of $M$ induced by any Morse function in the path component $\mathcal{N}$ of $M$. Moreover, Cerf’s proof of Proposition 4.2 shows that this instance of $\eta$ is a topological submersion and compatible with the respective stratifications, too.
We also give a brief description of what Cerf’s combinatorial description of strata of functions looks like in the discrete case.
Moreover, we consider the space of discrete Morse matchings, which has been introduced in [CJ05] and further investigated in [CM17] and [LS21]. We show that the space of discrete Morse matchings is also canonically associated to the Morse arrangement:
@@
Let $X$ be a CW complex, let $\mathcal{A}$ be the Morse arrangement on $\mathbb{R}^{X}$ and let $\mathfrak{M}(X)$ be the space of discrete Morse matchings on $X$. Then there is a canonical embedding of posets $\mathfrak{M}(X)\subset\mathcal{L}(\mathcal{A})$, where $\mathcal{L}(\mathcal{A})$ denotes the intersection poset of $\mathcal{A}$.
We describe the space of merge trees in a similar way as the space of discrete Morse functions. We model merge trees as maps from combinatorial merge trees to the real numbers and topologize their space using a mixture of the Euclidean distance and a kind of edit distance (see Definition 4.39). Our edit path distance is similar to an edit distance in [WG22], although their edit moves focus on edges, whereas our approach focuses on nodes.
@@
The edit path distance $\operatorname{d}$ from Definition 4.39 is a pseudo-metric on the space of merge trees $Mer$ and a metric on the space of strict merge trees $Mer_{<}$ and the space of well-branched merge trees $Mer_{wb}$.
@@
Let $X$ be a regular CW complex. Then the map $M\colon\mathcal{M}(X)\rightarrow Mer$ from Definition 4.26 that maps a discrete Morse function to its induced merge tree is Lipschitz continuous.
For the relationship to the space of barcodes, we use a similar approach [^1] One can identify the space of barcodes over a given combinatorial barcode in this work with the space $\mathbb{R}^{2n}$ in [BG23, Section 4.1]. as in [BG23] for the space of barcodes and a construction for the barcode induced by a merge tree similar to the one in [CDG+24]. We obtain the following result:
@@
The map $B\colon Mer_{wb}\rightarrow Bar$ from Definition 4.51 is Lipschitz continuous.
This way, we provide an alternative proof for the stability of merge trees and degree 0 barcodes, in our case with respect to notions of distance that are motivated by this Morse theoretic approach.
@@ Acknowledgements
The author would like to thank Max Planck Institute for Mathematics for the great scientific environment in which this project was started and the majority of the work was conducted. The author would also like to thank the Dioscuri Centre in TDA, located at the Mathematical Institute of the Polish Academy of Science, where the finishing touches were added. Moreover, the author would like to thank Andrea Bianchi, Florian Kranhold and Paul Mücksch for helpful discussions about the project. Most notably the author thanks his PhD advisor, Viktoriya Ozornova, for her advice, the many helpful discussions, and the detailed feedback at multiple occasions. Furthermore, the author thanks Clemens Bannwart for fruitful discussions on the topics of cell equivalences and gradient fields.
@@ 2. Hyperplane Arrangements and Their Associated Spaces
We review some well established notions from combinatorial geometry and topology, which will prove to be useful for our endeavor. The standard references we refer to are [BLS+99] and [AM17]. While most notions are standard in the literature, we adapt and extend some of them to the setting of this article.
Definition 2.1 (compare [BLS+99] and [AM17]).
@@
We call the path components of the complement $V\setminus\mathcal{A}$ the (open) regions of the arrangement $\mathcal{A}$. Moreover, we call intersections of half-spaces associated to $\mathcal{A}$, with at least one half-space per hyperplane chosen, faces of the hyperplane arrangement. For oriented hyperplane arrangements, each of these faces $\sigma$ is associated with a sign vector $w\in\{+,-,0\}^{\mathcal{A}}$, where ++ means that $\sigma$ lies on the positive side of the corresponding hyperplane, - means that $\sigma$ lies on the negative side of the corresponding hyperplane, and 0 means that $\sigma$ lies inside corresponding hyperplane. The regions correspond to sign vectors with no entry 0. The face poset $\Sigma(\mathcal{A})$ is the set of faces of $\mathcal{A}$ ordered by inclusion.
While hyperplane arrangements can be defined in arbitrary vector spaces, it is often useful to use coordinate spaces, i.e. spaces that come with a choice of basis. This allows us to use equations in the ordered coordinates in order to specify the arrangements we consider. We consider in particular partially ordered coordinate spaces.
@@
A partially ordered coordinate space is a real vector space $V$ together with a choice of a basis that is indexed over some partially ordered set $P$.
For convenience, we choose a specific categorical point of view for partially ordered coordinate spaces:
@@
The category of partially ordered coordinate spaces $po\operatorname{Vect}$ consists of
@@ 1. Objects:
    vector spaces $\operatorname{Map}(P,\mathbb{R})$ for finite posets $P$ together with the standard basis induced by $P$, and
@@ 2. Morphisms:
Remark 2.4.
The category of partially ordered coordinate spaces is related to the image of the contravariant functor $\operatorname{Map}(\_,\mathbb{R})\colon\operatorname{Poset}\rightarrow\operatorname{Vect}_{\mathbb{R}}$, where $\operatorname{Poset}$ denotes the category of finite posets and monotone maps, and $\operatorname{Vect}_{\mathbb{R}}$ denotes the category of finite-dimensional real vector spaces and linear maps. In addition to the data of the image of the functor, the specific choice of partially ordered basis given by the domain poset is part of data of partially ordered coordinate spaces.
In this work, we consider partially ordered coordinated spaces where the bases are indexed over either face posets of CW complexes or posets associated to merge trees or barcodes. We recall the definition of the face poset of any given CW complex.
@@
Let $X$ be a CW complex. We say that a cell $\sigma$ is a face of a cell $\tau$ if $\sigma\subseteq\tau$. The face poset of $X$ is the set of cells of $X$, denoted by $F(X)$, together with a partial order given by $\sigma\leq\tau$ if and only if $\sigma$ is a face of $\tau$.
We can characterize the three frameworks mentioned above as certain types of posets, which makes partially ordered coordinate vector spaces a unifying framework for them:
Definition 2.6.
A poset $P$ is of regular CW type if it is isomorphic to the face poset $F(X)$ of a regular CW complex $X$.
@@
    each non-minimal element $a\in P$ has at least two elements $b\neq b^{\prime}$ such that $b\prec a$ and $b^{\prime}\prec a$, where $\prec$ denotes cover relation. [^2] The cover relation $\prec$ in posets is defined as $a\prec b$ if and only if $a<b$, and there is no $c$ such that $a<c<b$.
@@
In order to see that Definition 2.6 models merge trees, we remark that (2) and (3) model arbitrary rooted trees: the chains in (3) correspond to shortest paths between the root and other nodes of the tree. The feature that inner nodes of merge trees correspond to mergers of path components, i.e. each inner node has at least two children, is reflected in (4).
@@
Example 2.8.
We consider the following example of a manifold with a height function, the induced merge tree and the corresponding poset of merge tree type.
[…]
Figure 1. An example of a merge tree induced by a height function on a manifold and the corresponding poset of merge tree type.
@@
We denote by $\mathcal{H}_{n}$ the hyperplane arrangement in $\mathbb{R}^{n}$ given by the equations $x_{i}=x_{j}$ for all $i\neq j$. We call $\mathcal{H}_{n}$ the braid arrangement $\mathcal{H}_{n}$ (see [BLS+99, Example 2.3.3]).
@@
If $\mathcal{A}$ is central, then $\mathcal{L}(\mathcal{A})$ is a geometric lattice of rank $r(\mathcal{A})$. Since all arrangements in this work are central, we also refer to $\mathcal{L}(\mathcal{A})$ as the intersection lattice of $\mathcal{A}$. It is straightforward to see that $\Pi(\mathcal{A})$ is never the face poset of a simplicial complex if $\mathcal{A}$ is central and $\lvert\mathcal{A}\rvert>1$.
On the other hand, the face poset $\Sigma(\mathcal{A})$ always has the structure of the face poset of a regular CW complex (see [AM17, 1.1.8]).
@@
We call a hyperplane arrangement simplicial if $\Sigma(\mathcal{A})\setminus\{\mathcal{Z}(\mathcal{A})\}$ has the structure of a face poset of a pure [^3] A simplicial complex is called pure if all its inclusion-maximal simplices have the same dimension $d$. simplicial complex.
In fact, $\Sigma(\mathcal{A})\setminus\{\mathcal{Z}(\mathcal{A})\}$ always gives a regular CW decomposition of the sphere in $\mathcal{Z}^{\perp}$. Hence, the real condition for a hyperplane arrangement being simplicial is that $\Sigma(\mathcal{A})\setminus\{\mathcal{Z}(\mathcal{A})\}$ needs to be isomorphic to the face poset of a simplicial complex. Then any such simplicial complex will be pure.
3. The Space of Discrete Morse Functions
Our goal is to present a framework which allows us to analyze geometric properties of, and homotopies between, discrete Morse functions on (regular) CW complexes. It turns out that manipulations of discrete Morse functions, like cancellations of critical cells and reordering critical cells by swapping their critical values can be geometrically realized by homotopies between discrete Morse functions, i.e. paths in the space of discrete Morse functions.
In order to define the space of discrete Morse functions on a CW complex, we consider the space of all discrete functions, which the space of discrete Morse functions will turn out to be a subspace of. As common when working with discrete Morse theory, we adopt a slight abuse of notation by not distinguishing between a CW complex and its set of cells.
@@ 3. The Space of Discrete Morse Functions
Let $X$ be a finite CW complex. We call the space $\mathbb{R}^{X}\coloneqq\mathbb{R}^{F(X)}\coloneqq\prod\limits_{\sigma\in X}\mathbb{R_{\sigma}}$ the space of discrete functions on $X$.
Here, we notice that the coordinates of the space of discrete functions are non-trivially partially ordered rather than just linearly ordered. This partial order will help us to investigate which of the discrete functions are discrete Morse functions.
@@
Let $X$ be a finite CW complex and let $F(X)$ be the face poset of $X$. A discrete Morse function on $X$ is a function $f\colon F(X)\rightarrow\mathbb{R}$ such that for every $p$-dimensional cell $\alpha^{(p)}\in X$
@@
    $\{\beta^{(p+1)}\supset\alpha|f(\beta)\leq f(\alpha)\}\leq 1$, and
@@
    $\{\gamma^{(p-1)}\subset\alpha|f(\gamma)\geq f(\alpha)\}\leq 1$.
@@
Cells for which the inequalities in (1) and (2) both are strict are called critical.
@@
We include the general version for non-regular CW complexes mostly for the comparison to the smooth case in Section 4.1. For obvious reasons, we can drop condition (3) if $X$ is regular. It is highly inconvenient to work with discrete Morse theory on non-regular CW complexes because in that case the face poset does not contain all information of the homotopy type at hand. Moreover, we make the definition for the non-regular case a bit stricter than Forman in order to simplify the definition of the induced gradient field, i.e. Morse functions are automatically strictly monotone along face inclusions of codimension higher than one.
We recall that by [FOR98, Lemma 2.5] for each cell $\alpha\in X$ at most one of the inequalities (1) and (2) from Definition 3.2 may actually be an equality, whereas the other one has to be a strict inequality. Therefore, for each cell $\alpha\in X$ such that (1)/(2) is an equality there is exactly one cell $\beta\supset\alpha$$\gamma\subset\alpha$ such that $f(\beta)\leq f(\alpha)$$f(\alpha)\leq f(\gamma)$. Hence, any discrete Morse function $f\colon F(X)\rightarrow\mathbb{R}$ induces a partial matching $\nabla f$ on $X$ by matching $\alpha$ with $u(\alpha)\coloneqq\beta$$d(\alpha)\coloneqq\gamma$. The unmatched cells are called critical. To be precise, $\nabla f$ is a partial matching on the Hasse diagram $D(X)$ of the face poset $F(X)$ of $X$:
@@
Let $X$ be a CW complex. The Hasse diagram of the face poset of $X$ is the directed graph $D(X)$ that has vertices $V(D(X))\coloneqq F(X)$, that is the cells of $X$, and an oriented edge $(\tau,\sigma)$ whenever $\sigma\subset\tau$ is a face relation such that there exists no $\theta\in X$ with $\sigma\subset\theta\subset\tau$ ($\sigma$ is covered by $\tau$).
@@
In case of a regular CW complex $X$, the directed edges correspond to face relations of codimension one. For non-regular CW complexes, edges might correspond to face relations of higher codimension. This simplifies explicit constructions of discrete Morse functions in the regular case because here, all necessary considerations can be checked exclusively on the face poset. In the more general case, the face relations that correspond to all intended matched pairs would need to be checked for regularity.
Furthermore, for non-regular CW complexes, the gradient paths needed for the boundary maps of the Morse chain complex are harder to spot inside the modified Hasse diagram, since they are then not just zig-zags between any matched edges and any unmatched egdges.
@@
Definition 3.6 ([BEN16, Section 2.1]).
Let $X$ be a CW complex. A Morse–Benedetti function on $X$ is a function on the face poset $f\colon F(X)\rightarrow\mathbb{R}$ that fulfills for all cells $\sigma,\tau$ in $X$ that:
  • Monotonicity:
    If $\sigma\subset\tau$ we have $f(\sigma)\leq f(\tau)$.
  • Semi-injectivity:
    $\lvert f^{-1}(\{z\})\rvert\leq 2$ for all $z\in\mathbb{R}$.
  • Genericity:
  • Regularity:
    If $\sigma\subset\tau$ is irregular, then $f(\sigma)<f(\tau)$.
For some purposes, it is useful to weaken Benedetti’s definition:
@@
Let $X$ be a CW complex. We call a function $f\colon F(X)\rightarrow\mathbb{R}$ a weak Morse–Benedetti function if it is a discrete Morse function and satisfies semi-injectivity and genericity in the sense of Definition 3.6.
Example 3.8.
We consider the following example of three discrete Morse function that induce the same matching.
[…]
Figure 2. Three different dMfs that induce the same matching (indicated in red). The one in the middle is weakly Morse–Benedetti but not Morse–Benedetti. The one on the right is Morse–Benedetti.
The dMf on the left is not weakly Morse–Benedetti due to the two critical edges labeled 5. The dMf in the middle is not Morse–Benedetti because of the matched pair with labels 2 and 3.
The definition of discrete Morse functions gives rise to a subdivision of the space of discrete functions induced by an oriented hyperplane arrangement in $\mathbb{R}^{X}$ as follows:
We observe that each point $p\in\mathbb{R}^{X}$ represents a discrete function $f\colon F(X)\rightarrow\mathbb{R}$ by $f(\sigma)=p_{\sigma}$. The conditions (1) and (2) (and (3)) of Definition 3.2 impose conditions on the components of a discrete function $p\in\mathbb{R}^{X}$ whenever there is a face relation of codimension 1 (or a cover relation) between the corresponding cells. That is, for the question whether $p$ represents a discrete Morse function it is relevant whether the inequality $p_{\sigma}<p_{\tau}$ holds whenever $\sigma\subset\tau$ is a face relation of codimension 1 (or a cover relation). This information is equivalent to the question on which side of the hyperplane $H_{\sigma}^{\tau}$ in $\mathbb{R}^{X}$ given by $x_{\sigma}=x_{\tau}$ the point $p$ lies. Hence, the conditions (1) and (2) (and (3)) of Definition 3.2 can be checked by analyzing the position of $p$ relative to the hyperplane arrangement $\mathcal{A}$ given by the equations $x_{\sigma}=x_{\tau}$ for all cover face relations [^4] Face relations of codimension 1 are a special case of cover relations. $\sigma\subset\tau$ in $X$. We define the following oriented hyperplane arrangement:
@@
Let $E$ be the set of cover face relations in $X$. We define the hyperplane arrangement $\mathcal{A}(X)\coloneqq\{H_{\sigma}^{\tau}|\sigma\subset\tau\in E\}$ where $H_{\sigma}^{\tau}$ is given by the equation $x_{\sigma}=x_{\tau}$. We observe that the assignment $(\sigma\subset\tau)\mapsto H_{\sigma}^{\tau}$ defines a bijection $E\cong\mathcal{A}$. Then let $\mathcal{S}(X)\coloneqq\{+,-,0\}^{E}$ be the sign vectors for the oriented hyperplane arrangement $\mathcal{A}^{+}(X)$ obtained by orienting $\mathcal{A}(X)$ in the following way: For each hyperplane $H_{\sigma}^{\tau}$ we define the open half-space defined by $x_{\sigma}<x_{\tau}$ to be the positive side of $H_{\sigma}^{\tau}$, the open half-space given by $x_{\sigma}>x_{\tau}$ to be the negative side of $H_{\sigma}^{\tau}$, and on $H_{\sigma}^{\tau}$ itself we have the value 0. We call $\mathcal{A}^{+}(X)$ the Morse arrangement on $\mathbb{R}^{X}$, respectively on $X$.
For any sign vector $x\in\mathcal{S}(X)$ we define the sets $x_{+}\coloneqq\{e\in E|x_{e}=+\}$, $x_{-}\coloneqq\{e\in E|x_{e}=-\}$, and $x_{0}\coloneqq\{e\in E|x_{e}=0\}$.